OpenStudy (anonymous):

Which lines are perpendicular to 5x + 4y = 8? Choose all answers that are correct. –5x + 4y = –5 5y = –4x – 15 5y = 4x + 5 –4x + 5y = 10

OpenStudy (anonymous):

-5x+4y=-5 and -4x + 5y=10 is perpendicular

OpenStudy (anonymous):

How? Can you please explain? :):):)

OpenStudy (igreen):

Perpendicular lines have the same slope.

OpenStudy (igreen):

Change 5x + 4y = 8 into the form of y = mx + b to find its slope. 5x + 4y = 8 Subtract 5x to both sides: 4y = -5x + 8 Divide 4 to all terms: y = -5/4x + 2 So the slope of 5x + 4y = 8 is -5/4.

OpenStudy (texaschic101):

perpendicular lines have negative reciprocal slopes. All that means is if you have a slope of 4/5....to find perpendicular, you " flip " the slope and change the sign...so the negative reciprocal of 4/5 is - 5/4. That being said...you will need to put the original equation in y = mx + b form, where m is your slope. 5x + 4y = 8 4y = -5x + 8 y = -5/4x + 2 but we need perpendicular...so we need the slope of 4/5...see how I flipped the slope and changed the sign. Now you would just have to go through all the answer choices, putting them in y = mx + b form....and the ones that have a slpe of 4/5 are your perpendicular lines. any questions ? Not too hard...just time consuming

OpenStudy (igreen):

Ohh, it said Perpendicular..xD

OpenStudy (igreen):

Perpendicular lines have OPPOSITE slope.

OpenStudy (texaschic101):

these aren't hard...just a little time consuming

OpenStudy (igreen):

5y = 4x + 5 --> y = 4/5x + 1 This is perpendicular to the original equation, 5x + 4y = 8.